OpenStudy (anonymous):

A force of 2i+7j N acts on a body of mass 5kg for 10 seconds. The body was initially moving with constant velocity of i−2j m/s. Find the final velocity of the body in m/s, in vector form. Having problem solving with vectors...

OpenStudy (amistre64):

isnt it just the adding of the two vectors?

OpenStudy (amistre64):

my first thought was dot product; but that doesnt result in a vector does it

OpenStudy (amistre64):

m1v1+m2v2 = (m1+m2)v3 .... is something that might apply

OpenStudy (anonymous):

F=ma a=(v_f-v_i)/t

OpenStudy (anonymous):

yes, the formula is F = m(v2-v1)/t but the vectors are my problem because the answer is to be given in the vector form

OpenStudy (amistre64):

shivams looks eloquentier to me

OpenStudy (amistre64):

10F/m +v1 = v2

OpenStudy (anonymous):

@amistre64 Ft/m = v_f - v_i

OpenStudy (anonymous):

@amistre64 , I wanted to say is this new word-eloquentier ??

OpenStudy (amistre64):

10<2,7> -------- + <1,-2> = Vf 2

OpenStudy (amistre64):

lol, might be new; if so i might as well patent it :)

OpenStudy (anonymous):

See. vector subraction A vector = x i + y j +z k B vector = p i + q j + rk B-A = (p-x) i + (q-y) j + (r -z) k

OpenStudy (amistre64):

i never did care much for the ijk format

OpenStudy (amistre64):

how does the N play into the picture? that Newton?

OpenStudy (anonymous):

@amistre64 , @Prof.Grace is new to this. let him get used to this. Later he can easily learn what you are trying to convey :D

OpenStudy (anonymous):

F is vector and newton is unit of F

OpenStudy (anonymous):

@Prof.Grace , any problem now with vector addition and subtraction??

OpenStudy (amistre64):

should Vi be converted to a Netwon equivalent?

OpenStudy (anonymous):

No. Because Ft/M is m/s

OpenStudy (amistre64):

hmm, looks good to me then. Math i can handle, trying to physicate ... not so much :)

OpenStudy (anonymous):

@amistre64 , Maths and physics are always inter-related :D

OpenStudy (anonymous):

i think i've gotten the answer, thanks y'all..

OpenStudy (anonymous):

Good luck :D

OpenStudy (anonymous):

If you were given a constant acceleration and the time and also the initial velocity, you could find final velocity by v=u+at. You can do the exactly same thing for vectors as well as I’ve explained below. Divide the vector force F with the scalar mass to get the accelaration vector. Multiply this with time to get the change in velocity vector. Add this vector to the original velocity vector. And you've got the final velocity vector. acc vector a= (2i+7j)/5 m s-2 change in velocity δv= (2i+7j)/5 m s-2 * 10s =4i+14j m s-1 Add this to the original velocity vector.